# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 4.3. The Crandall–Ishii–Lions Theorem on Sums
    the function $u$ is said to be $\mathcal{F}$-subharmonic on $X$ if $u$ is $\mathcal{F}$-subharmonic at each $x\in X$.
    The space of all $\mathcal{F}$-subharmonic functions on $X$ will be denoted by $\mathcal{F}(X)$.
    The corresponding (and standard) notion of $u\in\operatorname{USC}(X)$ being a viscosity subsolution of the PDE (LABEL:Fxududdu=0) in $x\in X$ is that
@@
    The analogous notion of $u\in\operatorname{LSC}(X)$ being a viscosity supersolution of the PDE (LABEL:Fxududdu=0) in $x\in X$ is that
@@
    In the case that a constraint $\mathcal{F}\subset\mathcal{J}^{2}(X)$ is determined by a differential operator $F\in C(\mathcal{J}^{2}(X))$ by (LABEL:compatibility), the definition of supersolution (5.8) suggests defining $u\in\operatorname{LSC}(X)$ to be $\mathcal{F}$-superharmonic in $x\in X$ if
@@
    Finally, a function $u\in C(X)$ is said to be $\mathcal{F}$-harmonic in $x\in X$ if both (5.6) and (5.9) hold. In the next section, we will use a natural notion of duality to reduce superharmonics to subharmonics of a dual constraint set.
@@
    In order to be meaningful, the viscosity formulation of $\mathcal{F}$-subharmonic functions of Definition LABEL:defn:FSH requires making a monotonicity assumption on the constraint set $\mathcal{F}$ with respect to the $\mathcal{S}(n)$ variable. More precisely one needs to assume that a constraint set $\mathcal{F}$ fiberwise satisfies the positivity property:
    for $\text{for each $x\in X$:}\quad(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P\geqslant 0\ \text{in}\ \mathcal{S}(n).$ (5.10)
@@
    Let $u\in\mathcal{F}(X)$ with $u(x_{0})\neq-\infty$ and suppose that $\varphi$ is an upper test function for $u$ at $x_{0}$. Since $u$ is $\mathcal{F}$-subharmonic in $x_{0}$ we have $(r,p,A):=J^{2}_{x_{0}}\varphi\in\mathcal{F}_{x_{0}}$ but the function
    $\psi:=\varphi+\frac{1}{2}\langle P_{0}(x-x_{0}),x-x_{0}\rangle$
    is also an upper test function for $u$ at $x_{0}$ but, by (5.11), satisfies
@@
    =$\mathcal{F}=\{(x,r,p,A)\in\mathcal{J}^{2}(X):\ F(x,r,p,A)\geqslant 0\},$ (5.12)
    positivity of $\mathcal{F}$ is implied by $F$ being degenerate ellipitc
    $(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P>0,$ (5.13)
    while positivity of $\mathcal{F}$ implies that $F$ satisfies $F(x,r,p,A+P)\geqslant 0$ for all $P\geqslant 0$ if $F(x,r,p,A)\geqslant 0$.
    This monotonicity ((5.10) or (5.13)) should be viewed as the weakest possible notion of ellipticity suitable for a meaningful viscosity theory in both contexts.
@@
    Of fundamental importance in many constructions is to have the following coherence property for $\mathcal{F}$-subharmonics in the sense of Definition LABEL:defn:FSH with $\mathcal{F}\neq\emptyset$: if $u$ is $C^{2}$ near $x_{0}\in X$ then
    $u$ is $\mathcal{F}$-subharmonic at $x_{0}\ \Leftrightarrow\ J^{2}_{x_{0}}u\in\mathcal{F}_{x_{0}}$. (5.14)
    The implication ()($\Rightarrow$)( ⇒ ) is obvious since $u$ is its own upper test function in $x_{0}$. The reverve implication ()($\Leftarrow$)( ⇐ ) requires the positivity condition (5.10). Indeed, suppose that $J^{2}_{x_{0}}u\in\mathcal{F}_{x_{0}}$ and that $\varphi$ is $C^{2}$ upper test function for $u$ in $x_{0}$. Since $u-\varphi$ has a local maximum value of 00 in $x_{0}$, by elementary calculus, one has
    $\varphi(x_{0})=u(x_{0}),D\varphi(x_{0})=Du(x_{0}),\text{and}\ D^{2}\varphi(x_{0})\geqslant Du(x_{0}),$
    and denoting by $P:=D^{2}\varphi(x_{0})-Du(x_{0})\geqslant 0$ in $\mathcal{S}(n)$ and hence
@@
    Remark 5.4 (Equivalent notions of $\mathcal{F}$-subharmonicity).
    In the notation of Remark 2.8, Definition LABEL:defn:FSH says that
    $u$ is $\mathcal{F}$-subharmonic in $x\ \Leftrightarrow\ (u(x),p,A)\subset\mathcal{F}_{x}\quad\forall\,(p,A)\in J_{C^{2}}(x,u)$, (5.15)
    where $J_{C^{2}}(x,u)$ is the space of $C^{2}$ upper contact jets for $u$ in $x$ ($(p,A)=(D\varphi(x),D^{2}\varphi(x)$ with $\varphi$ a $C^{2}$ upper contact function for $u$ in $x$). As noted in Remark 2.8, in the chain of inclusions (2.8)
    [8, Lemma C.1] shows that all of these spaces have the same closure in $\mathcal{J}^{2}$ for each $x\in X$. Hence, if
    the fiber x is closed in $\mbox{the fiber $\mathcal{F}_{x}$ is closed in $\mathcal{J}^{2}$},$ (5.17)
    then in (5.15) one can replace the space of upper contact jets $J_{C^{2}}(x,u)$ with any of the others in the chain (5.16). Moreover, if
    $\mathcal{F}$ is closed in $\mathcal{J}^{2}(X)$, (5.18)
    then one can also replace $J_{C^{2}}(x,u)$ with a limiting space $\overline{J}_{C^{2}}(x,u)$ consisting of those $(p,A)\in{\mathbb{R}}^{n}\times\mathcal{S}(n)$ for which there are sequences $\{x_{k}\}_{k\in\mathbb{N}}\subset X$ and $(p_{k},A_{k}\}_{k\in\mathbb{N}}\subset J_{C^{2}}(x,u)$ for which
    $(x_{k},u(x_{k}),p_{k},A_{k})\to(x,u(x),p,A))\quad\text{as $k\to+\infty$},$
    and simarly for the closures of the other spaces of upper test jets.
    The considerations above lead us to a first class of constraint sets $\mathcal{F}$ which have a meaningful and flexible notion of upper semicontinuous $\mathcal{F}$-subharmonic functions.
@@
    and $\mathcal{F}$ satisfies fiberwise the positivity condition:
    for $\text{for each $x\in X$:}\quad(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P>0.$ (P)
    As noted in Remarks 5.2 and 5.3, the positivity condition (P) is essential for a meaningful notion of upper semicontinuous $\mathcal{F}$-subharmonics and in order to have $C^{2}$-coherence, while the closedness condition (C) ensures the equivalent formulations of Remark 5.4. Condition (C) also ensures that the elementary properties of $\mathcal{F}(X)$ involving limits and upper envelopes hold (see [24, Theorem 2.6]; the same properties are also listed, for nonconstant coefficient subequations, in Proposition 7.4).
    The two minimal conditions (C) and (P), however, are not sufficient for all of our purposes. First, as already mentioned, in order to treat the natural Dirichlet problem for $\mathcal{F}$-harmonic functions, we will need to prove comparison; that is, to show that
@@
    whenever $u$ and $v$ are $\mathcal{F}$-subharmonic and $\mathcal{F}$-superharmonic respectively (or a sub/super-solution pair for a degenerate elliptic PDE). In order to avoid obvious counterexamples to comparison (5.19), one needs to require that $\mathcal{F}$ has another fiberwise monotonicity property with respect to ${\mathbb{R}}$ variable: that is, that $\mathcal{F}$ satisfies the negativity property
    for each $x\in X:\quad(r,p,A)\in\mathcal{F}_{x}\implies(r+s,p,A)\in\mathcal{F}_{x}\quad\forall s<0$. (N)
@@
    Property (N) will have additional implications and utility, including the fact that it guarantees a sliding property for subharmonics: if $u$ is a subharmonic, then so is $u-m$ for each constant $m\geqslant 0$. Indeed, if $(p,A)$ is an upper contact jet for $u$ at $x$, so it will be for $u-m$. This sliding property will play a key role in some upcoming constructions and is included, along with many elementary properties mentioned above, in Proposition 7.4. Hence, in general, one also wants property (N) for a constraint set $\mathcal{F}$. There are important exceptions to this, including when treating generalized principal eigenvalues and associated eigendirections where typically one has the opposite monotonicity in the $r$ variable (see [6], for example.)
@@
    is implied by the operator $F$ being proper elliptic [^9] In the standard viscosity theory, $F$ is often said to be merely proper. We prefer the term proper elliptic to recall the crucial importance of weak (degenerate) ellipticity required for the viscosity theory; that is, for each $(x,r,p,A)\in\mathcal{J}^{2}(X)$
    $F(x,r+s,p,A)+P)\geqslant F(x,r,p,A)\quad\forall\,s\leqslant 0\ \text{in}\ {\mathbb{R}},P\geqslant 0\ \text{in}\ \mathcal{S}(n).$ (5.20)
@@
    Note that if some fiber $\mathcal{F}_{x}$ of a constraint set were to be empty, then $X$ might not admit an $\mathcal{F}$-subharmonic. To see this, the same argument we used at the beginning of the proof of Proposition 3.19 shows that the space of upper quadratic test jets $J^{2,+}_{x}u$ for $u$ at $x$ is non empty for some $x\in X$ and hence for each $(p,A)\in J^{2,+}_{x}u$ the condition $(u(x),p,A)\in\mathcal{F}_{x}$ fails if $\mathcal{F}_{x}$ is empty.
    Moreover, if some fiber $\mathcal{F}_{x}$ were to be equal to the entire space of $2$-jets $\mathcal{J}^{2}$, the property of being $\mathcal{F}$-subharmonic in $x$ is satisfied by all functions with are upper semicontinuous in some neighborhood of $x$. Hence one should consider only constraint sets $\mathcal{F}$ such that
    for each $x\in X:\quad\mathcal{F}_{x}$ is closed, non empty with $\mathcal{F}_{x}\neq\mathcal{J}^{2}$.
    Finally, one obtains a well behaved and useful theory of constraint sets $\widetilde{\mathcal{F}}$ which are dual to a given constraint set $\mathcal{F}$ provided that one also assumes the three conditions of topological stability
    $\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}},\qquad(\operatorname{int}{\mathcal{F}})_{x}=\operatorname{int}(\mathcal{F}_{x}),\ \forall\,x\in X\qquad\mathcal{F}_{x}=\overline{\operatorname{int}{\mathcal{F}}_{x}},\ \forall\,x\in X.$ (T)
@@
    A subequation (constraint set) on an open set $X\subset{\mathbb{R}}^{n}$ is subset $\mathcal{F}\subset$ with nontrivial fibers (that is, proper and nonempty fibers) which satisfies the positivity property
    for $\text{for each $x\in X$:}\quad(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P>0.$ (P)
@@
    for each $x\in X:\quad(r,p,A)\in\mathcal{F}_{x}\implies(r+s,p,A)\in\mathcal{F}_{x}\quad\forall s<0$. (N)
@@
    $\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}},\qquad(\operatorname{int}{\mathcal{F}})_{x}=\operatorname{int}(\mathcal{F}_{x}),\ \forall\,x\in X\qquad\mathcal{F}_{x}=\overline{\operatorname{int}{\mathcal{F}}_{x}},\ \forall\,x\in X.$ (T)
    In addition to ensuring a well-behaved duality including the reflexivity property $\widetilde{\widetilde{\mathcal{F}}}=\mathcal{F}$, property (T) also ensures the local existence of strict classical subharmonics, when we adopt the following notion of strictness.
@@
    strictly $\mathcal{F}$-subharmonic on $X$ if this inclusion holds for each $x\in X$.
@@
    Using the topological condition (T), one can prove that for each $x\in X$ there exists a quadratic function which is strictly $\mathcal{F}$-subharmonic on a neighborhood of $x$. Indeed, by (T) we know that $\operatorname{int}{\mathcal{F}}_{x}\neq\emptyset$ for all $x\in X$; hence, let $J=(r,p,A)\in\operatorname{int}{\mathcal{F}}_{x}$ and consider the quadratic function $\varphi_{J}$ associated to $J$, namely
@@
    Two interesting remarks that one finds in [26] are worth noting. They play an important role in the Almost Everywhere Theorem (cf. Theorem 8.1).
@@
    The notion of being $\mathcal{F}$-subharmonic can be expressed in terms of an inclusion between subequations. Indeed, each upper semicontinuous function $w$ on $X$ determines a minimal primitive subequation $\mathcal{G}$ such that $w$ is $\mathcal{G}$-subharmonic, namely $\mathcal{G}=\overline{\mathcal{J}^{2,+}w}$, where
    $\mathcal{J}^{2,+}w:=\{(x,w(x),p,A):\ x\in X,(p,A)\in J^{2,+}_{x}w\}$ (5.21)
    Therefore $w$ is $\mathcal{F}$-subharmonic if and only if $\mathcal{J}^{2,+}w\subset\mathcal{F}$, if and only if $\mathcal{G}\subset\mathcal{F}$ since $\mathcal{F}$ is closed.
@@
    Note that we already found a subset which surely contains $\mathcal{J}^{2,+}w$. Indeed, if we set
@@
    then Lemma 2.15 says that if $w$ is locally quasi-convex and $E$ has full measure, then $\mathcal{J}^{2,+}w\subset\overline{\mathcal{J}(w,E)}$.
@@
    A (primitive) subequation $\mathcal{F}$ is said a constant coefficient, or translation invariant, (primitive) subequation if the mapping $x\mapsto\mathcal{F}_{x}$ is constant, that is if the fibers do not depend on the point $x$.
@@
    Since our definition of subequation requires that all the fibers are nontrivial, it is important to specify the open set $X$ on which we are considering a subequation $\mathcal{F}$. Nevertheless, if $\mathcal{F}$ is a constant coefficient subequation on $X$, it can be naturally extended to the whole ${\mathbb{R}}^{n}$ by setting, for some $x_{0}\in X$ fixed, $\mathcal{F}_{y}=\mathcal{F}_{x_{0}}$ for all $y\in{\mathbb{R}}^{n}$. Hence any constant coefficient subequation can be thought of as defined on all of ${\mathbb{R}}^{n}$ and hence identify them with a nontrivial subset of $\mathcal{J}^{2}\vcentcolon={\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ that satisfies the conditions (P), (N) and (T)). Notice that in this case, the property (T) reduces to the single condition $\mathcal{F}=\overline{\operatorname{int}\mathcal{F}}$ in $\mathcal{J}^{2}$. Obviously, when speaking about $\mathcal{F}$-subharmonics/superharmonics/harmonics, we must specify on which open set $X$ we are working.
    Finally, we present a few representative examples of constant coefficient subequations on an arbitrary open set $X$ in ${\mathbb{R}}^{n}$, starting from the classical case of (Laplacian) subharmonics.

### v3

@@ 5.1. Subequation constraint sets and their subharmonics
[…]
The corresponding (and standard) notion of $u\in\operatorname{USC}(X)$ being a viscosity subsolution of the PDE (5.1) in $x\in X$ is that
@@
The analogous notion of $u\in\operatorname{LSC}(X)$ being a viscosity supersolution of the PDE (5.1) in $x\in X$ is that
@@
[…]
@@
[…]
@@
[…]
$(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P\geqslant 0\ \text{in}\ \mathcal{S}(n),\quad\forall x\in X.$ (5.9)
@@
[…]
$\psi\vcentcolon=\varphi+\frac{1}{2}\langle P_{0}(x-x_{0}),x-x_{0}\rangle$
is also an upper test function for $u$ at $x_{0}$, but, by (5.10), it satisfies
@@
positivity of $\mathcal{F}$ is implied by $F$ being degenerate ellipitc, i.e.
$F(x,r,p,A)\leqslant F(x,r,p,A+P)\quad\forall P\geqslant 0,$
while positivity of $\mathcal{F}$ implies that $F$ satisfies
$F(x,r,p,A)\geqslant 0\quad\implies\quad F(x,r,p,A+P)\geqslant 0\quad\forall P\geqslant 0.$
This monotonicity should be viewed as the weakest possible notion of ellipticity suitable for a meaningful viscosity theory in both contexts.
@@
[…]
The implication ()($\Rightarrow$)( ⇒ ) is obvious since $u$ is its own upper test function at $x_{0}$. The reverse implication ()($\Leftarrow$)( ⇐ ) requires the positivity condition (5.9). Indeed, suppose that $J^{2}_{x_{0}}u\in\mathcal{F}_{x_{0}}$ and that $\varphi$ is $C^{2}$ upper test function for $u$ at $x_{0}$. Since $u-\varphi$ has a local maximum value of 00 in $x_{0}$, by elementary calculus, one has
$\varphi(x_{0})=u(x_{0}),\quad D\varphi(x_{0})=Du(x_{0}),\quad\text{and}\quad D^{2}\varphi(x_{0})\geqslant Du(x_{0}),$
hence, denoting by $P\vcentcolon=D^{2}\varphi(x_{0})-Du(x_{0})\geqslant 0$ in $\mathcal{S}(n)$,
@@
[…]
In the notation of Remark 2.8, Definition 5.1 says that
[…]
As noted in Remark 2.8, we have the chain of inclusions (2.6) and [9, Lemma C.1] shows that all of those spaces have the same closure in $\mathcal{J}^{2}$ for each $x\in X$. Hence, if the fiber $\mathcal{F}_{x}$ is closed in $\mathcal{J}^{2}$, then in (5.12) one can replace the space of upper contact jets $J_{C^{2}}(x,u)$ with any of the others in the chain (2.6). Moreover, if $\mathcal{F}$ is closed in $\mathcal{J}^{2}(X)$, then one can also replace $J_{C^{2}}(x,u)$ with a limiting space $\overline{J}_{C^{2}}(x,u)$ consisting of those $(p,A)\in{\mathbb{R}}^{n}\times\mathcal{S}(n)$ for which there are sequences $\{x_{k}\}_{k\in\mathbb{N}}\subset X$ and $\{(p_{k},A_{k})\}_{k\in\mathbb{N}}\subset J_{C^{2}}(x,u)$ for which $(x_{k},u(x_{k}),p_{k},A_{k})\to(x,u(x),p,A)$ as $k\to\infty$, and similarly for the other spaces of upper test jets.
[…]
@@
and $\mathcal{F}$ satisfies fiberwise the positivity condition
$(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P>0,\quad\forall x\in X.$ (P)
[…]
@@
[…]
$(r,p,A)\in\mathcal{F}_{x}\implies(r+s,p,A)\in\mathcal{F}_{x}\quad\forall s<0,\quad\forall x\in X.$ (N)
@@
Property (N) will have additional implications and utility, including the fact that it guarantees a sliding property for subharmonics: if $u$ is a subharmonic, then so is $u-m$ for each constant $m\geqslant 0$. Indeed, if $(p,A)$ is an upper contact jet for $u$ at $x$, so it will be for $u-m$. This sliding property will play a key role in some upcoming constructions and is included, along with many elementary properties mentioned above, in Proposition 7.5. Hence, in general, one also wants property (N) for a constraint set $\mathcal{F}$. However, there are important exceptions to this, including when treating generalized principal eigenvalues and associated eigendirections where typically one has the opposite monotonicity in the $r$ variable (see [6], for example).
@@
is implied by the operator $F$ being proper elliptic [^2] In the standard viscosity theory, $F$ is often said to be merely proper. We prefer the term proper elliptic to recall the crucial importance of weak (degenerate) ellipticity required for the viscosity theory.; that is, for each $(x,r,p,A)\in\mathcal{J}^{2}(X)$,
$F(x,r+s,p,A+P)\geqslant F(x,r,p,A)\quad\forall\,s\leqslant 0\ \text{in}\ {\mathbb{R}},\ \ P\geqslant 0\ \text{in}\ \mathcal{S}(n).$
@@
[…]
Finally, one obtains a well behaved and useful theory of constraint sets $\widetilde{\mathcal{F}}$ which are dual to a given constraint set $\mathcal{F}$ provided that one also assumes the following three conditions of topological stability:
$\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}},\qquad(\operatorname{int}{\mathcal{F}})_{x}=\operatorname{int}(\mathcal{F}_{x})\quad\forall x\in X,\qquad\mathcal{F}_{x}=\overline{\operatorname{int}{\mathcal{F}}_{x}}\quad\forall x\in X.$ (T)
@@
A subequation (constraint set) on an open set $X\subset{\mathbb{R}}^{n}$ is a subset $\mathcal{F}\subset\mathcal{J}^{2}(X)$ with nontrivial fibers (that is, proper and non-empty fibers) which satisfies the positivity property
$(r,p,A)\in\mathcal{F}_{x}\implies(r,p,A+P)\in\mathcal{F}_{x}\quad\forall P>0,\quad\forall x\in X,$ (P)
@@
$(r,p,A)\in\mathcal{F}_{x}\implies(r+s,p,A)\in\mathcal{F}_{x}\quad\forall s<0,\quad\forall x\in X,$ (N)
@@
$\mathcal{F}=\overline{\operatorname{int}{\mathcal{F}}},\qquad(\operatorname{int}{\mathcal{F}})_{x}=\operatorname{int}(\mathcal{F}_{x})\quad\forall x\in X\qquad\mathcal{F}_{x}=\overline{\operatorname{int}{\mathcal{F}}_{x}}\quad\forall x\in X.$ (T)
In addition to ensuring a well-behaved duality including the reflexivity property (see Section 6.1) property (T) also ensures the local existence of strict classical subharmonics, when we adopt the following notion of strictness.
@@
[…]
@@
[…]
@@
Two interesting remarks that one finds in [29] are worth noting. They play an important role in the Almost Everywhere Theorem (Theorem 8.1).
@@
[…]
$\mathcal{J}^{2,+}w\vcentcolon=\{(x,w(x),p,A):\ x\in X,\ (p,A)\in J^{2,+}_{x}w\}$ (5.14)
[…]
@@
We already found a subset which surely contains $\mathcal{J}^{2,+}w$. Indeed, if we set
@@
then Lemma 2.15 says that if $w$ is locally semiconvex and $E$ has full measure, then $\mathcal{J}^{2,+}w\subset\overline{\mathcal{J}(w,E)}$.
@@
A (primitive) subequation $\mathcal{F}$ is said a constant coefficient, or translation invariant, (primitive) subequation if the mapping $x\mapsto\mathcal{F}_{x}$ is constant; that is if the fibers do not depend on the point $x$.
@@
[…]
5.2. Some simple classical examples
Finally, we present a few representative examples of constant coefficient subequations on an arbitrary open set $X\subset{\mathbb{R}}^{n}$, starting from the classical case of Laplacian subharmonics.